# SAM 2: Segment Anything in Images and Videos

### v1

@@ 1 Introduction
We are releasing our work under permissive open licences, including the SA-V dataset (CC by 4.0) a version of the model SAM 2 (Apache 2.0), along with an interactive online demo at https://sam2.metademolab.com.
@@ 2 Related work
Segment Anything (Kirillov et al., 2023) introduces a promptable image segmentation task where the goal is to output a valid segmentation mask given an input prompt such as a bounding box or a point that refers to the object of interest. SAM trained on the SA-1B dataset allows for zero-shot segmentation with flexible prompting, which enabled its adoption to a wide range of downstream applications. Recent work has extended SAM by improving its quality. For example, HQ-SAM (Ke et al., 2024) enhances SAM by introducing a High-Quality output token and training the model on fine-grained masks. Another line of work focuses on SAM’s efficiency to enable wider use in real-world and mobile applications, such as EfficientSAM (Xiong et al., 2023), MobileSAM (Zhang et al., 2023a), and FastSAM (Zhao et al., 2023). The success of SAM led to its adoption in a wide range of applications, such as medical imaging (Ma et al., 2024; Deng et al., 2023; Mazurowski et al., 2023; Wu et al., 2023a), remote sensing (Chen et al., 2024; Ren et al., 2024), motion segmentation (Xie et al., 2024), and camouflaged object detection (Tang et al., 2023).
@@
Interactive video object segmentation has emerged as a crucial task to efficiently obtain object segmentations in videos (masklets) with user guidance, often in the form of scribbles, clicks, or bounding boxes. A few early approaches (Wang et al., 2005; Bai & Sapiro, 2007; Fan et al., 2015) deploy graph-based optimization to guide the segmentation annotation process. More recent approaches (Heo et al., 2020; Cheng et al., 2021b; Delatolas et al., 2024) often adopt a modular design, converting the user inputs into a mask representation on a single frame and then propagating it to other frames. Our work shares a similar goal to these works to segment objects across videos with a good interactive experience, and we build a strong model along with a large and diverse dataset in pursuit of this goal.
In particular, the DAVIS interactive benchmark (Caelles et al., 2018) allows interactively segmenting an object via scribble inputs on multiple frames. Inspired by the DAVIS interactive benchmark, we also adopt an interactive evaluation setting for the promptable video segmentation task in §6.1.
@@
Click-based input is easier to collect (Homayounfar et al., 2021) for interactive video segmentation. Recent works have used a combination of SAM on images with video trackers based on masks (Cheng et al., 2023b; Yang et al., 2023; Cheng et al., 2023c) or points (Rajič et al., 2023). However, these approaches have limitations: the tracker may not work for all objects, SAM may not perform well for image frames from videos, and there is no mechanism to interactively refine a model’s mistakes, other than re-annotating using SAM from scratch on the erroneous frame and restarting the tracking from there.
Semi-supervised Video Object Segmentation (VOS).
@@
Semi-supervised VOS usually begins with an object mask as input in the first frame, which must be accurately tracked throughout the video (Pont-Tuset et al., 2017). It is called “semi-supervised” since the input mask can be seen as a supervision signal of the object appearance that is available only for the first frame. This task has drawn significant attention due to its relevance in various applications, including video editing, robotics, and automatic background removal.
@@
Early neural network-based approaches have often used online fine-tuning on the first video frame (Caelles et al., 2016; Perazzi et al., 2016; Yoon et al., 2017; Maninis et al., 2017; Hu et al., 2018a; Bhat et al., 2020; Robinson et al., 2020) or on all frames (Voigtlaender & Leibe, 2017) to adapt the model to the target object. Faster inference has been achieved with offline-trained models, conditioned either only on the first frame (Hu et al., 2018b; Chen et al., 2018), or also integrating the previous frame (Oh et al., 2018; Yang et al., 2018, 2020). This multi-conditioning has been extended to all frames with RNNs (Xu et al., 2018a) and cross-attention (Oh et al., 2019; Cheng et al., 2021a; Li et al., 2022a; Yang et al., 2021b, 2024; Cheng & Schwing, 2022; Yang & Yang, 2022; Wang et al., 2022; Cheng et al., 2023a; Goyal et al., 2023). Recent approaches (Zhang et al., 2023b; Wu et al., 2023b) extend a single vision transformer to jointly process the current frame along with all previous frames and associated predictions, resulting in a simple architecture but at a prohibitive inference cost. Semi-supervised VOS can be seen as a special case of our Promptable Visual Segmentation (PVS) task, as it is equivalent to only providing a mask prompt in the first video frame. Nevertheless, annotating the required high-quality object mask in the first frame is practically challenging and time-consuming.
@@
Many datasets have been proposed to support the VOS task. Early VOS datasets (Prest et al., 2012; Li et al., 2013; Ochs et al., 2014; Fan et al., 2015), such as DAVIS (Pont-Tuset et al., 2017; Caelles et al., 2019), include high-quality annotations but their limited size does not allow training deep-learning based approaches. Covering 94 object categories over 4 thousand videos, YouTube-VOS (Xu et al., 2018b) is the first large-scale dataset for the VOS task. As algorithms became better and benchmark performance started to saturate, researchers have looked at increasing the difficulty of the VOS task by specifically focusing on occlusions (Qi et al., 2022; Ding et al., 2023), long videos (Hong et al., 2023, 2024), extreme transformations (Tokmakov et al., 2022), object diversity (Wang et al., 2021b, 2023) or scene diversity (Athar et al., 2022).
@@ 3 Task: promptable visual segmentation
The PVS task allows providing prompts to the model on any frame of a video. Prompts can be positive/negative clicks, bounding boxes, or masks, either to define an object to segment or to refine a model-predicted one. To provide an interactive experience, upon receiving a prompt on a specific frame, the model should immediately respond with a valid segmentation mask of the object on this frame. After receiving initial (one or multiple) prompts (either on the same frame or different frames), the model should propagate these prompts to obtain the masklet of the object across the entire video, which contains the segmentation mask of the target object on every video frame. Additional prompts can be provided to the model on any frame to refine the segment throughout the video (example in Fig. 2). For details on the task, see §A.
SAM 2, introduced in the next section (§4), is applied as a data collection tool to the PVS task for building our SA-V dataset (§5). The model is evaluated (§6) in an online and offline setting by simulating interactive video segmentation scenarios involving annotations across multiple frames, in the conventional semi-supervised VOS setting where annotations are limited to the first frame, and for image segmentation on the SA benchmarks.
@@ 4 Model
Our model can be seen as a generalization of SAM to the video (and image) domain. SAM 2 (Fig. 3) supports point, box, and mask prompts on individual frames to define the spatial extent of the object to be segmented across the video. For image input, the model behaves similarly to SAM. A promptable and light-weight mask decoder accepts a frame embedding and prompts (if any) on the current frame and outputs a segmentation mask for the frame. Prompts can be iteratively added on a frame in order to refine the masks.
Unlike SAM, the frame embedding used by the SAM 2 decoder is not directly from an image encoder and is instead conditioned on memories of past predictions and prompted frames. It is possible for prompted frames to also come “from the future” relative to the current frame. Memories of frames are created by the memory encoder based on the current prediction and placed in a memory bank for use in subsequent frames. The memory attention operation takes the per-frame embedding from the image encoder and conditions it on the memory bank to produce an embedding that is then passed to the mask decoder.
We describe individual components and training below and provide more details in Appendix C.
@@
Our prompt encoder is identical to SAM’s and can be prompted by clicks (positive or negative), bounding boxes, or masks to define the extent of the object in a given frame. Sparse prompts are represented by positional encodings summed with learned embeddings for each prompt type, while masks are embedded using convolutions and summed with the frame embedding.
@@
Unlike SAM where there is always a valid object to segment given a positive prompt, in the PVS task it is possible for no valid object to exist on some frames (e.g. due to occlusion). To account for this new output mode, we add an additional head that predicts whether the object of interest is present on the current frame. Another difference from SAM is that we use skip connections from the hierarchical image encoder (bypassing the memory attention) to incorporate high-resolution information for mask decoding (see §C).
@@
In addition to the spatial memory, we store a list of object pointers as lightweight vectors for high-level semantic information of the object to segment, based on mask decoder output tokens of each frame (Meinhardt et al., 2022). Our memory attention cross-attends to both spatial memory features and these object pointers.
@@
The model is trained jointly on image and video data. Similar to previous work (Kirillov et al., 2023; Sofiiuk et al., 2022), we simulate interactive prompting of the model. We sample sequences of 8 frames and randomly select up to 2 frames to prompt and probabilistically receive corrective clicks which are sampled using the ground-truth masklet and model predictions during training. The training task is to sequentially (and “interactively”) predict the ground-truth masklet. Initial prompts to the model can be the ground-truth mask with probability $0.5$, a positive click sampled from the ground-truth mask with probability $0.25$, or a bounding box input with probability $0.25$. See §C for more details.
@@ 5.1 Data engine
Despite an improvement in annotation time, this decoupled approach requires annotating masks in intermediate frames from scratch, without previous memory. We then advanced to develop the fully-featured SAM 2, capable of performing both interactive image segmentation and mask propagation in a unified model.
@@
Table 2: Segmentation accuracy ($\mathcal{J}\&\mathcal{F}$ metric) improvement from adding data from each data engine phase. “VOS” is a set of video object segmentation datasets. Details are in §E.
In Table 2, we show the performance comparison of SAM 2 trained on the available data at the end of each phase keeping the number of iterations fixed, therefore measuring solely the impact of the additional data. We evaluate on our own SA-V val set and also on 9 zero-shot benchmarks (see §E.1 for details) using the standard $\mathcal{J}\&\mathcal{F}$ accuracy metric (the higher the better) when prompting with 3-clicks on the first frame. We note a consistent improvement after iteratively including the data from each phase, not only on the in-domain SA-V val set, but also on the 9 zero-shot benchmarks.
@@ 5.2 SA-V dataset
We collected a new set of 50.9K videos captured by crowdworkers. Videos comprise 54% indoor and 46% outdoor scenes with an average duration of 14 seconds. Videos feature “in-the-wild” diverse environments, and cover various everyday scenarios. Our dataset has more videos than existing VOS datasets, and as shown in Fig. 5, videos span 47 countries and were captured by diverse participants (self-reported demographics).
@@
See Appendix D for more details on the data engine and SA-V dataset.
@@
(a) Size
(b) Geography
Male274Female236
i Gender
18-2410925-4030541-6488
$i$ Age
(c) Crowdworker Demographics
@@ 6 Zero-shot experiments
Here, we compare SAM 2 with previous work on zero-shot video tasks (§6.1) and image tasks (§6.2). We report the standard $\mathcal{J}\&\mathcal{F}$ metric (Pont-Tuset et al., 2017) for video and mIoU metric for image tasks. Unless otherwise mentioned, the results reported in this section follow our default setup using Hiera-B+ image encoder with a resolution of 1024 and trained on the full combination of datasets, i.e., SAM 2 (Hiera-B+) in Table 7 (see also §C.2 for details).
6.1 Video tasks
6.1.1 Promptable video segmentation
@@
We first evaluate promptable video segmentation, which involves simulating an interactive setting that resembles the user experience. We have two settings, offline evaluation, where multiple passes are made through a video to select frames to interact with based on the largest model error, and online evaluation, where the frames are annotated in a single forward pass through the video. These evaluations are conducted on 9 densely annotated zero-shot video datasets using $N_{\mathrm{click}}=3$ clicks per frame (see §E.1 for details).
We create two strong baselines, SAM+XMem++ and SAM+Cutie, based on two state-of-the-art models for video object segmentation, XMem++ (Bekuzarov et al., 2023) and Cutie (Cheng et al., 2023a). We use XMem++ to generate a video segmentation based on mask inputs on one or multiple frames. SAM is used to provide an initial mask or to refine an output (by feeding the current segmentation as a mask prompt to SAM). For the SAM+Cutie baseline, we modify Cutie to allow taking mask inputs on multiple frames.
@@
6.1.2 Semi-supervised video object segmentation
@@
SAM 2 64.3 73.2 75.4 72.9 77.6
Table 4: Zero-shot accuracy across 17 video datasets under semi-supervised VOS evaluation using different prompts. The table shows the averaged $\mathcal{J}\&\mathcal{F}$ for each type of prompt (1, 3 or 5 clicks, bounding boxes, or ground-truth masks) in the first video frame (: in this case we directly use masks as inputs into XMem++ or Cutie without using SAM).
We next evaluate the semi-supervised video object segmentation (VOS) setting (Pont-Tuset et al., 2017) with click, box, or mask prompts only on the first frame of the video. When using click prompts, we interactively sample either 1, 3 or 5 clicks on the first video frame, and then track the object based on these clicks.
Similar to the interactive setting in §6.1.1, we compare to XMem++ and Cutie, using SAM for click and box prompts, and in their default setting when using mask prompts. We report the standard $\mathcal{J}\&\mathcal{F}$ accuracy (Pont-Tuset et al., 2017), except for on VOST (Tokmakov et al., 2022), where we report the $\mathcal{J}$ metric following its protocol. The results are in Table 4. SAM 2 outperforms both baselines on the 17 datasets, using various input prompts. The results underline that SAM 2 also excels at the conventional, non-interactive VOS task with mask input, for which these other works are specifically designed. More details are in §E.1.3.
6.1.3 Fairness evaluation
1-click 3-click mask
gender
male 81.9 95.1 95.9
female 75.1 94.1 95.2
age
18-26 77.2 95.0 95.7
26-50 76.7 94.7 95.8
50+ 81.4 95.1 96.2
In Appendix G, we provide model, data and annotation cards for SA-V.
6.2 Image tasks

### v2

@@ 1 Introduction
We are releasing our work under permissive open licences, including the SA-V dataset (CC by 4.0), the SAM 2 model checkpoints [^1] All the results presented in this paper are based on a new version of SAM 2 (improved over our initial release; denoted as “SAM 2.1” in https://github.com/facebookresearch/sam2), which we will refer to as SAM 2 throughout for brevity., training code (Apache 2.0), and code for our interactive online demo (Apache 2.0).
@@ 2 Related work
Segment Anything (Kirillov et al., 2023) introduces a promptable image segmentation task where the goal is to output a valid segmentation mask given an input prompt such as a bounding box or a point that refers to the object of interest. SAM trained on the SA-1B dataset allows for zero-shot segmentation which enabled its adoption to a wide range of applications. Recent work has extended SAM, e.g., by introducing a High-Quality output token to train on fine-grained masks (Ke et al., 2024), or improve SAM’s efficiency (Xiong et al., 2023; Zhang et al., 2023a; Zhao et al., 2023). More broadly, SAM is used in a wide range of applications, including medical imaging (Ma et al., 2024; Deng et al., 2023; Mazurowski et al., 2023; Wu et al., 2023a), remote sensing (Chen et al., 2024; Ren et al., 2024), motion segmentation (Xie et al., 2024), and camouflaged object detection (Tang et al., 2023).
@@
Interactive video object segmentation has emerged as a crucial task to efficiently obtain object segmentations in videos (masklets) with user guidance, often in the form of scribbles, clicks, or bounding boxes. A few early approaches (Wang et al., 2005; Bai & Sapiro, 2007; Fan et al., 2015) deploy graph-based optimization to guide the segmentation annotation process. More recent approaches (Heo et al., 2020; Cheng et al., 2021b; Delatolas et al., 2024) often adopt a modular design, converting user inputs into a mask representation on a single frame and then propagating it to other frames.
@@
Click-based input is easier to collect (Homayounfar et al., 2021) for interactive video segmentation. Recent works have used a combination of SAM on images with video trackers based on masks (Cheng et al., 2023b; Yang et al., 2023; Cheng et al., 2023c) or points (Rajič et al., 2023). However, these approaches have limitations: the tracker may not work for all objects, SAM may not perform well on video frames, and there is no mechanism to interactively refine a model’s mistakes, other than re-annotating using SAM in each frame and restarting the tracking from there.
Our work shares a similar goal to these works to segment objects across videos interactively, and we build a strong unified model that directly takes prompts for interactive video segmentation, along with a large and diverse dataset in pursuit of solving this goal.
Video Object Segmentation (VOS).
@@
The VOS task begins with an object mask as input in the first frame, which must be accurately tracked throughout the video (Pont-Tuset et al., 2017). The task is referred to as “semi-supervised VOS” since the input mask can be seen as supervision signal of the object which is available only in the first frame. This task has drawn significant attention due to its relevance in applications, including video editing or robotics.
@@
Early deep learning based approaches have often used online fine-tuning on the first video frame (Caelles et al., 2016; Perazzi et al., 2016; Yoon et al., 2017; Maninis et al., 2017; Hu et al., 2018a; Bhat et al., 2020; Robinson et al., 2020) or on all frames (Voigtlaender & Leibe, 2017) to adapt the model to the target object. Faster inference has been achieved with offline-trained models, conditioned either only on the first frame (Hu et al., 2018b; Chen et al., 2018), or also integrating the previous frame (Oh et al., 2018; Yang et al., 2018, 2020). This multi-conditioning has been extended to all frames with RNNs (Xu et al., 2018a) and transformers (Oh et al., 2019; Cheng et al., 2021a; Li et al., 2022a; Yang et al., 2021b, 2024; Cheng & Schwing, 2022; Yang & Yang, 2022; Wang et al., 2022; Cheng et al., 2023a; Goyal et al., 2023; Zhang et al., 2023b; Wu et al., 2023b).
Semi-supervised VOS can be seen as a special case of our Promptable Visual Segmentation (PVS) task, with only a mask prompt in the first video frame. Notably, annotating the required high-quality object mask in the first frame in VOS is practically challenging and time-consuming for inference.
@@
Many datasets have been proposed to support the VOS task. Early VOS datasets (Prest et al., 2012; Li et al., 2013; Ochs et al., 2014; Fan et al., 2015), such as DAVIS (Pont-Tuset et al., 2017; Caelles et al., 2019), include high-quality annotations but their size limits deep-learning based approaches. YouTube-VOS (Xu et al., 2018b) is the first large-scale dataset for VOS. As algorithms became better and benchmark performance started to saturate, researchers have looked at increasing the difficulty of the VOS task by specifically focusing on occlusions (Qi et al., 2022; Ding et al., 2023), long videos (Hong et al., 2023, 2024), extreme transformations (Tokmakov et al., 2022), object diversity (Wang et al., 2021b, 2023) or scene diversity (Athar et al., 2022).
@@ 3 Task: promptable visual segmentation
Our PVS task allows providing prompts to the model on any frame of a video. Prompts can be positive/negative clicks, boxes, or masks, either to define an object to segment or to refine a model-predicted one. To provide an interactive experience, upon receiving a prompt on a specific frame, the model should immediately respond with a valid segmentation mask of the object on this frame. After receiving initial prompts (either on the same frame or different frames), the model should propagate these prompts to obtain the masklet of the object across the entire video, localizing the segmentation mask of the target on every video frame. Additional prompts can be provided to the model on any frame to refine the segment throughout the video (example in Fig. 2). For details on the task, see §B.
SAM 2 (§4) is applied as a data collection tool to the PVS task for building our SA-V dataset (§5). We evaluate the model (§6) by simulating interactive video segmentation scenarios across multiple frames, in the conventional semi-supervised VOS setting where annotations are limited to the first frame, and for image segmentation on the SA benchmarks.
@@ 4 Model
SAM 2 (Fig. 3) can be seen as a generalization of SAM to the video (and image) domain, taking point, box, and mask prompts on individual frames to define the spatial extent of the object to be segmented spatio-temporally. Spatially, the model behaves similarly to SAM. A promptable and light-weight mask decoder takes an image embedding and prompts (if any) and outputs a segmentation mask for the frame. Prompts can be iteratively added on a frame in order to refine the masks.
The frame embedding used by the SAM 2 decoder is not directly from an image encoder and is instead conditioned on memories of past predictions and prompted frames. It is possible for prompted frames to also come “from the future” relative to the current frame. Memories of frames are created by the memory encoder based on the current prediction and placed in a memory bank for use in subsequent frames. The memory attention operation takes the per-frame embedding from the image encoder and conditions it on the memory bank, before the mask decoder ingests it to form a prediction.
We describe individual components and training below and provide more details in Appendix D.
@@
Our prompt encoder is identical to SAM’s and can be prompted by clicks (positive or negative), boxes, or masks to define the extent of the object in a given frame. Sparse prompts are represented by positional encodings summed with learned embeddings for each prompt type, while masks are embedded using convolutions and summed with the frame embedding.
@@
Unlike SAM where there is always a valid object to segment given a positive prompt, in the PVS task it is possible for no valid object to exist on some frames (e.g. due to occlusion). To support this new output mode, we add an additional head that predicts whether the object of interest is present on the current frame. Another novelty are skip connections from our hierarchical image encoder (bypassing the memory attention) to incorporate high-resolution embeddings for mask decoding (see §D).
@@
In addition to the spatial memory, we store a list of object pointers as lightweight vectors for high-level semantic information of the object to segment, based on mask decoder output tokens of each frame. Our memory attention cross-attends to both spatial memory features and these object pointers.
@@
The model is trained jointly on image and video data. Similar to previous work (Kirillov et al., 2023; Sofiiuk et al., 2022), we simulate interactive prompting of the model. We sample sequences of 8 frames and randomly select up to 2 frames to prompt and probabilistically receive corrective clicks which are sampled using the ground-truth masklet and model predictions during training. The training task is to sequentially (and “interactively”) predict the ground-truth masklet. Initial prompts to the model can be the ground-truth mask with probability $0.5$, a positive click sampled from the ground-truth mask with probability $0.25$, or a bounding box input with probability $0.25$. See §D for more details.
@@ 5.1 Data engine
Despite an improvement in annotation time, this approach requires annotating masks in intermediate frames from scratch without previous memory. We then advanced to develop the fully-featured SAM 2, capable of both interactive segmentation and mask propagation in a unified model.
@@
Table 2: Segmentation accuracy ($\mathcal{J}\&\mathcal{F}$ metric) improvement from adding data from each data engine phase. “VOS” is a set of video object segmentation datasets. Details are in §F.
In Table 2, we show the performance comparison of SAM 2 trained on the available data at the end of each phase keeping the number of iterations fixed, therefore measuring solely the impact of the additional data. We evaluate on our own SA-V val set and also on 9 zero-shot benchmarks (see §F.1 for details) using the standard $\mathcal{J}\&\mathcal{F}$ accuracy metric (the higher the better) when prompting with 3-clicks on the first frame. We note a consistent improvement after iteratively including the data from each phase, not only on the in-domain SA-V val set, but also on the 9 zero-shot benchmarks.
@@ 5.2 SA-V dataset
We collected a new set of 50.9K videos captured by crowdworkers. Videos comprise 54% indoor and 46% outdoor scenes with an average duration of 14 seconds. Videos feature “in-the-wild” diverse environments, and cover various everyday scenarios.
@@
See Appendix E for more details on the data engine and SA-V dataset, including a fairness evaluation.
@@ 6 Zero-shot experiments
Here, we compare SAM 2 with previous work on zero-shot video and image tasks. We report the standard $\mathcal{J}\&\mathcal{F}$ metric (Pont-Tuset et al., 2017) for video and mIoU metric for image tasks. Unless otherwise mentioned, the results in this section follow our default setup using Hiera-B+ image encoder with a resolution of 1024 and trained on the full combination of datasets, i.e., SAM 2 (Hiera-B+) in Table 6 (see also §D.2 for details).
6.1 Promptable video segmentation
@@
We first evaluate promptable video segmentation, which involves simulating an interactive setting that resembles the user experience. We have two settings, offline evaluation, where multiple passes are made through a video to select frames to interact with based on the largest model error, and online evaluation, where the frames are annotated in a single forward pass through the video. These evaluations are conducted on 9 densely annotated zero-shot video datasets using $N_{\mathrm{click}}=3$ clicks per frame (see §F.1 for details).
We create two strong baselines, SAM+XMem++ and SAM+Cutie, based on two state-of-the-art models for video object segmentation, XMem++ (Bekuzarov et al., 2023) and Cutie (Cheng et al., 2023a).
We use XMem++ to generate a video segmentation based on mask inputs on one or multiple frames. SAM is used to provide an initial mask or to refine an output (by feeding the current segmentation as a mask prompt to SAM). For the SAM+Cutie baseline, we modify Cutie to allow taking mask inputs on multiple frames.
@@
6.2 Semi-supervised video object segmentation
@@
SAM 2 64.7 75.3 77.6 74.4 79.3
Table 4: Zero-shot accuracy across 17 video datasets using different prompts. We report average accuracy for each type of prompt (1, 3 or 5 clicks, bounding boxes, or ground-truth masks) in the first video frame (: this case directly uses masks as inputs into XMem++ or Cutie without SAM).
We evaluate the semi-supervised video object segmentation (VOS) setting (Pont-Tuset et al., 2017) with click, box, or mask prompts only on the first frame of the video. When using click prompts, we interactively sample either 1, 3 or 5 clicks on the first video frame.
Similar to the interactive setting in §6.1, we compare to XMem++ and Cutie, using SAM for click and box prompts, and in their default setting when using mask prompts. We report the standard $\mathcal{J}\&\mathcal{F}$ accuracy (Pont-Tuset et al., 2017), except for on VOST (Tokmakov et al., 2022), where we report the $\mathcal{J}$ metric following its protocol. The results are in Table 4. SAM 2 outperforms both methods on the 17 datasets. The results underline that SAM 2 also excels at the conventional, non-interactive VOS task with mask input, for which these other works are specifically designed. Details are in §F.1.3.
6.3 Image segmentation